# Re-Introducing LayerNorm: Geometric Meaning, Irreversibility and a Comparative Study with RMSNorm

### v1

@@ Abstract
Layer normalization is a pivotal step in the transformer architecture. This paper delves into the less explored geometric implications of this process, examining how LayerNorm influences the norm and orientation of hidden vectors in the representation space. We show that the definition of LayerNorm is innately linked to the uniform vector, defined as $\bm{1}=[1,1,1,1,\cdots,1]^{T}\in\mathbb{R}^{d}$. We then show that the standardization step in LayerNorm can be understood in three simple steps: (i) remove the component of a vector along the uniform vector, (ii) normalize the remaining vector, and (iii) scale the resultant vector by $\sqrt{d}$, where $d$ is the dimensionality of the representation space. We also introduce the property of ”irreversibility” for LayerNorm, where we show that the information lost during the normalization process cannot be recovered. In other words, unlike batch normalization, LayerNorm cannot learn an identity transform. While we present possible arguments for removing the component along the uniform vector, the choice of removing this component seems arbitrary and not well motivated by the original authors. To evaluate the usefulness of this step, we compare the hidden representations of LayerNorm-based LLMs with models trained using RMSNorm and show that all LLMs naturally align representations orthogonal to the uniform vector, presenting the first mechanistic evidence that removing the component along the uniform vector in LayerNorm is a redundant step. Our findings support the use of RMSNorm over LayerNorm as it is not only more computationally efficient with comparable downstream performance, but also learns a similar distribution of hidden representations that operate orthogonal to the uniform vector.
@@ 1 Introduction
The transformer architecture (Vaswani et al., 2017) has been the cornerstone of most recent advances in artificial intelligence and has rapidly become the architecture of choice in natural language processing, computer vision, speech, and various other domains. Layer normalization, a crucial yet often overlooked component of the transformer architecture, plays an integral role in stabilizing the training process in transformer-based models. Introduced by Ba et al. (2016), layer normalization standardizes the features at each layer, adjusting and scaling the activations to have zero mean and unit variance within a vector. This normalization is performed independently for each hidden vector in contrast to batch normalization (Ioffe & Szegedy, 2015), which relies on statistics (mean and variance) from a batch of data points. Layer normalization is especially effective when working with long sequences of variable lengths. While the benefits of layer normalization, such as improved training speed and better convergence, are well documented (Ba et al., 2016; Xu et al., 2019; Zhang & Sennrich, 2019), its specific effects on the hidden vectors within a model and the global properties of the resulting representations remain surprisingly underexplored.
In this paper, we first discuss the global effects of layer normalization on a vector. The conventional explanation of the LayerNorm operation is usually as follows: standardize each vector by subtracting the mean of its elements, divide by the standard deviation. While this is an accurate procedural definition, we ask a more global question: How does the LayerNorm operation transform a vector in representation space? We present a novel interpretation of LayerNorm and show that it can be understood in three steps: (i) throw away the component of the given vector along $\bm{1}=[1,1,1,1,\cdots,1]^{T}\in\mathbb{R}^{d}$, (ii) normalize the resultant vector; and (iii) scale the resulting vector by $\sqrt{d}$, where $d$ is the dimensionality of the vector space. This process is illustrated in Figure 1. This operation throws away the information along $\bm{1}=[1,1,1,1,\cdots 1]^{T}$, which we call the uniform vector, indicating that the information along the uniform vector may not be important.
We then present the property of “irreversibility,” where we show that the information lost in the LayerNorm process cannot be recovered. While “reversibility” is an established property in BatchNorm (Ioffe & Szegedy, 2015), this property has not been discussed in literature for layer normalization to the best of our knowledge. These results naturally lead to discussion about the importance of information along the uniform vector, which is lost during LayerNorm irreversibly, and its comparison with RMSNorm (Zhang & Sennrich, 2019), a variant of LayerNorm that doesn’t remove the component along the uniform vector and is used to train the latest Llama models (Touvron et al., 2023a; b; Dubey et al., 2024). Figure 2(a) shows the difference between the two normalization methods. We empirically show that RMSNorm-based models naturally produce hidden representations orthogonal to the uniform vector, thus rendering the removal of the uniform vector in LayerNorm inconsequential in practice.
@@
(b) This figure shows the projection of the original vector along the uniform vector, $(\bm{x}.\hat{\bm{1}})\hat{\bm{1}}$, and the remaining component, $\bm{x^{\bm{1}}_{\perp}}$.
(c) The component of the original vector, after removing the projection along the uniform vector, is kept ( $\bm{x^{\bm{1}}_{\perp}}$, shown in yellow).
(d) Normalizing perpendicular component to the uniform vector ($\bm{x^{\bm{1}}_{\perp}}$) to unit norm.
@@
    Provide a simple geometric interpretation of the layer normalization process that provides an intuitive explanation of how LayerNorm transforms a vector in representation space.
@@
    Introduce the property of “irreversibility” for LayerNorm and show that information lost in the layer normalization process cannot be recovered.
@@
    Show that removing the component along the uniform vector is a redundant step in LayerNorm, and models trained using RMSNorm also learn to create representations orthogonal to the uniform vector. These findings support using RMSNorm over LayerNorm.
  4. 4.
    Finally, we also present LayerNorm in action with experiments depicting its importance in stabilizing the residual stream vectors. LayerNorm not only regulates the ever-increasing norm of hidden vectors inside LLMs but also rotates each vector to orient it properly within representation space.
2 Background
@@
In this paper, we refer to the combination of step-1 and step-2 as the ”standardization step”, whereas step-3 is referred to as the ”scale-and-shift” step.
2.1 Computations in Decoder-only LLMs
In this paper, we study the hidden representations of modern decoder-only large language models (LLMs). Let $h^{l}$ represent the intermediate hidden state vectors between each decoder layer. As depicted in Figure 2(b), the computations within a layer of most decoder-only LLMs proceed as follows:
[…]
$a^{l}$ $=\texttt{Att}(f^{l})$ (2)
[…]
$m^{l}$ $=W^{l}_{proj}\sigma(W^{l}_{fc}g^{l}+b^{l}_{fc})+b_{proj}$ (4)
$h^{l}$ $=h^{l-1}+a^{l}+m^{l}$ (5)
(a) A diagrammatic explanation of LayerNorm and RMSNorm.
(b) A high-level diagram representing the computation within one decoder block of an LLM.
Figure 2: Combined figures showing the diagrams for LayerNorm and RMSNorm alongside the computation within a decoder block of an LLM.
$h^{l}=h^{0}+\sum^{i=l}_{i=0}+a^{i}+\sum^{i=l}_{i=0}+m^{i}$ (6)
3 Re-Introducing Layer Normalization
While the original definition of LayerNorm (section 2) gives us an operational description, it tells us very little about the global properties of the resulting vectors. For example, with the steps described in section 2, we get very little insight about the norm or the orientation of the standardized vector ($\bm{y}$). All we know is that the components of the vector $\bm{y}$ are standardized to have zero mean and unit variance within the vector. But how is this standardized vector oriented compared to the original vector and what is its norms? Such an explanation was also absent in the original formulation of layer normalization (Ba et al., 2016).
Initial explorations into the geometry of layer normalization by Brody et al. (2023) demonstrated that $\bm{y}$ maintains a norm of $\sqrt{d}$ and is oriented orthogonal to the constant vector $\bm{1}=[1,1,1,1,\cdots,1]^{T}\in\mathbb{R}^{d}$, where $d$ represents the dimensionality of the vector space of hidden vectors. Furthermore, a popular variant of LayerNorm, known as RMSNorm (Zhang & Sennrich, 2019), which omits the “mean subtraction” in the standardization step, has come out as a viable alternative to LayerNorm. In this paper, we expand upon these foundational studies by presenting a novel interpretation of the global effects of LayerNorm on a vector, which we argue provides a more intuitive and informative description of how LayerNorm modifies a hidden vector in representation space.
@@ 3 Re-Introducing Layer Normalization
$\mu=\frac{1}{d}\bm{1}^{T}\bm{x}$ (7)
where $\bm{1}=[1,1,1,1,\cdots 1]^{T}$ such that $\bm{1}\in\mathbb{R}^{d}$, which we refer to as the uniform vector in this paper. Using the scalar mean value calculated above, we define the mean vector as $\bm{\mu}=\mu\bm{1}$, where $\mu$ is a scalar mean as calculated in equation 7. Thus, the mean vector is a vector with each component equal to the mean of the components of vector $\bm{x}$. Using this, we can also rewrite the standard deviation of the components of $\bm{x}$ as:
@@ 3.1 The Uniform Vector and the Mean Vector
The uniform vector or $\bm{1}=[1,1,1,1,\cdots 1]^{T}$, is called so because all its components are equal or uniform and set to 1. This vector plays a very important role in the formulation and understanding of layer normalization. An important thing to note here is the norm of the uniform vector: $\|\bm{1}\|_{2}=\sqrt{\bm{1}^{T}\bm{1}}=\sqrt{d}$.
The mean vector in the context of LayerNorm is defined in a non-traditional manner. Traditionally, a mean vector is the sum of a few vectors. But in our definition, the mean vector, $\bm{\mu}=\mu\bm{1}$, is a scaled version of the uniform vector. It is scaled by a value that is the mean or the average of all the components of the hidden vector and is oriented in the direction of the uniform vector. Note that the mean vector is a function of an underlying vector $\bm{x}$, as is implicit in the calculation of $\mu$ in equation 7. An interesting property of the mean vector is its norm. Let’s define $\theta_{x1}$ as the angle between vectors $\bm{x}$ and the uniform vector. Then:
$\begin{split}\|\bm{\mu}\|&=\|\mu\bm{1}\|=\mu\|\bm{1}\|\\ &=\Big{(}\frac{1}{d}\bm{1}^{T}\bm{x}\Big{)}\|\bm{1}\|\\ &=\frac{1}{d}\|\bm{1}\|^{2}\|\bm{x}\|\cos\theta_{x1}=\|\bm{x}\|\cos\theta_{x1}\end{split}$ (10)
Here we replace the formula for the mean from equation 7, expand the inner product between the uniform vector and $\bm{x}$ in terms of their norms and the angle between them, and use the fact that $\|\bm{1}\|_{2}=\sqrt{d}$.
The norm of the mean vector gives us very important insights about what’s going on in the layer normalization process. Equation 10 shows that the norm of the mean vector is nothing but the projection of the underlying vector $\bm{x}$ along the uniform vector. By definition, the mean vector, $\bm{\mu}=\mu\bm{1}$, is oriented along the direction of the uniform vector. In other words, given that $\theta$ is the angle between vectors $\bm{x}$ and the uniform vector,
$\begin{split}\|\bm{\mu}\|=\|\bm{x}\|\cos\theta_{x1}=\bm{x}.\hat{\bm{1}}\end{split}$ (11)
@@
To the best of our knowledge, the geometrical meaning of the mean vector has not been discussed in literature before our work. In the next section, we incorporate this information to explain the LayerNorm process, giving a very intuitive and informative description of the process.
@@ 3.2 Explanation of Layer Normalization
If we go through the steps of layer normalization as discussed in section 2, the component-wise subtraction of the mean of all components of a vector can now be written completely in terms of the vector $\bm{x}$. We define $\bm{x^{\bm{1}}_{\perp}}$ as the component of $\bm{x}$ orthogonal to the uniform vector as follows:
$\bm{x^{\bm{1}}_{\perp}}=\bm{x}-\bm{\mu}=\bm{x}-(\bm{x}.\hat{\bm{1}})\hat{\bm{1}}$ (13)
@@
$\bm{y}=\sqrt{d}\frac{\bm{x^{\bm{1}}_{\perp}}}{\|\bm{x^{\bm{1}}_{\perp}}\|_{2}}$ (14)
where $d$ is the dimensionality of the vector being normalized. This shows that layer normalization can be simply defined as the normalization of the component of a vector orthogonal to the uniform vector, accompanied by a scaling factor. Figure 1 shows an intuitive description of this process. With this, we present the simple and elegant definition of the layer normalization process with a deep geometric meaning. The most intuitive recipe of the standardization step in LayerNorm is as follows:
The Standardization Step in LayerNorm
1:Throw away the component of a vector along the uniform vector, $\vec{\bm{1}}=[1,1,1,1,\cdots 1]^{T}$ (Fig 1(c))
2:Normalize the remaining vector (Fig 1(d))
3:Scale the resulting vector by $\sqrt{d}$, where $d$ is the dimensionality of the vector space (Fig 1(e))
Equation 14 also shows that the norm of the standardized vector is $\sqrt{d}$ and is oriented orthogonal to the direction of the uniform vector, which means it exists in a $d-1$ dimensional subspace orthogonal to the uniform vector (Brody et al., 2023).
@@ 3.3 The Irreversibility of Layer Normalization
The idea of reversibility is discussed briefly while introducing batch normalization (Ioffe & Szegedy, 2015), which normalizes each feature dimension in the input vector $\bm{x}$ independently by calculating the mean and standard deviation statistics over a large sample of such vectors. Specifically, let $\bm{x}^{1},\bm{x}^{2},\dots,\bm{x}^{b}$ be a set of $b$ vectors in the training set, each vector of dimension $d$. Then, both the batch normalization and layer normalization processes for each component can be represented in the following two-step process:
@@
The key difference between batch normalization and layer normalization lies in how the expectation values and variance are calculated for each component. In the case of batch normalization, the mean and variance for each dimension are calculated over the training set, i.e.
$\mathbb{E}[x_{i}]$ $=\frac{1}{b}\sum^{j=b}_{j=1}x^{j}_{i}$ (17)
$Var(x_{i})$ $=\frac{1}{b}\sum^{j=b}_{j=1}(x^{j}_{i}-\mathbb{E}[x^{j}_{i}])^{2}$ (18)
@@
Note that for batch normalization, $\mathbb{E}[x_{i}]$ and $Var(x_{i})$ are the same for each vector $\bm{x}^{j}\in\{\bm{x}^{1},\bm{x}^{2},\dots,\bm{x}^{b}\}$, but different for each component of the vector. In total, for a $d$-dimensional hidden vector space, there are $2d$ statistics in batch normalization, $d$ for the mean and $d$ for variance. This means that the information lost during the standardization step in batch normalization, if important, can be recovered in the scale-and-shift step by simply learning $\alpha_{i}=\sqrt{Var(x_{i})}$ and $\beta_{i}=\mathbb{E}[x_{i}]$, since we have $2d$ learnable parameters in batch normalization. The scale-and-shift step was a very conscious design choice of the inventors of batch normalization to allow for batch normalization to represent an identity transformation if the original hidden representations were optimal for the network.
A subtle but very important difference with layer normalization is how the expectation and variances are calculated for each component. For layer normalization, these statistics are calculated for each vector independently. Thus,
$\mathbb{E}[x_{i}]$ $=\mathbb{E}[x^{j}_{i}]=\frac{1}{d}\sum^{i=d}_{i=1}x^{j}_{i}$ (19)
$Var(x_{i})$ $=Var(x^{j}_{i})\frac{1}{d}\sum^{i=b}_{i=1}(x^{j}_{i}-\mathbb{E}[x^{j}_{i}])^{2}$ (20)
@@
We recommend the reader carefully note the summation indexes and normalization values above and compare them against batch normalization. The above equations show that $\mathbb{E}[x_{i}]$ and $Var(x_{i})$ in layer normalization are the same for each component $x_{i}$ within a vector but are different for different vectors. This means that two variables are created in LayerNorm per vector. As LLMs will normalize way more than $2d$ vectors throughout the course of their training, the number of statistic variables used in LayerNorm is of the order of the number of tokens used to train the model. Thus, the information lost during layer normalization cannot be recovered by learning just $2d$ components of the $\bm{\alpha}$ and $\bm{\beta}$ learnable parameters. This shows that information once lost during layer normalization cannot be recovered during the scale-and-shift step, making the layer normalization process irreversible. In other words, the learnable parameters in LayerNorm cannot represent an identity transformation. To the best of our knowledge, the “irreversibility” of layer normalization has not been discussed previously in literature.
3.4 What’s so Special about the Uniform Vector?
We see that the definition of LayerNorm is innately linked with the uniform vector, possibly unintended by the original authors. Since the information along the uniform vector is being removed irreversibly during layer normalization (section 3.3), the layer normalization process implicitly assumes that the information along the uniform vector is either not important or that the model should not store information along that direction.
To understand what removing the component along the uniform vector does, we need to go back to the component-wise definition of LayerNorm. Removing the component along the uniform vector is the same as subtracting the mean of components of a vector from each component. If all components of a vector are equal, this operation reduces the vector to a vector of zeros. Thus, LayerNorm forces the model to learn hidden representations such that all components of the vector are not the same; otherwise, the vector is reduced to zero. A possible argument for enforcing this could be that if all components of a vector are the same, the vector loses expressive power. This is a possible conjecture offered by us to potentially justify such a choice. If this is indeed the explanation, the need for such a step should be justified by showing that the hidden representations within a language model have a propensity towards having all components equal to each other.
A different perspective to understand the implications of the “mean subtraction” step can be seen using the global picture of LayerNorm presented in this paper. Two randomly chosen vectors in high dimensional spaces are nearly orthogonal to each other with a very small spread. Since the “mean subtraction” step in LayerNorm means removing the component along the arbitrarily chosen uniform vector, the need for such a step would be justified if the hidden representations created by the model had significant components along this vector. The choice of removing information along the uniform vector seems like an unintended choice without much justification given in literature, including the original paper Ba et al. (2016). RMSNorm, a popular variant of LayerNorm used to train the Llama series of models, does not perform this step and performs just as well. This provides a strong motivation to explore the need for the “mean subtraction” step in LayerNorm (section 4.3).
Table 1: List of models used in experiments.
Model Name Model Dim (d) Num Params Num Layers Norm. Type

### v2

# Geometric Interpretation of Layer Normalization and a Comparative Analysis with RMSNorm

@@ Abstract
This paper presents a novel geometric interpretation of LayerNorm and explores how LayerNorm influences the norm and orientation of hidden vectors in the representation space. With these geometric insights, we prepare the foundation for comparing LayerNorm with RMSNorm. We show that the definition of LayerNorm is innately linked to the uniform vector, defined as $\boldsymbol{1}=[1,1,1,1,\cdots,1]^{T}\in\mathbb{R}^{d}$. We then show that the standardization step in LayerNorm can be understood in three simple steps: (i) remove the component of a vector along the uniform vector, (ii) normalize the remaining vector, and (iii) scale the resultant vector by $\sqrt{d}$, where $d$ is the dimensionality of the representation space. We also provide additional insights into how LayerNorm operates at inference time. Finally, we compare the hidden representations of LayerNorm-based LLMs with models trained using RMSNorm and show that all LLMs naturally operate orthogonal to the uniform vector at inference time, that is, on average they do not have a component along the uniform vector during inference. This presents the first mechanistic evidence that removing the component along the uniform vector in LayerNorm is a redundant step. These results advocate for using RMSNorm over LayerNorm which is also more computationally efficient.
Geometric Interpretation of Layer Normalization and
a Comparative Analysis with RMSNorm
Akshat Gupta, Atahan Ozdemir, Gopala Anumanchipalli UC Berkeley {akshat.gupta, atahanozdemir, gopala}@berkeley.edu
Figure 1: A diagrammatic explanation of LayerNorm and RMSNorm.
@@ 1 Introduction
The transformer architecture (Vaswani et al., 2017) has been the cornerstone of most recent advances in artificial intelligence and has rapidly become the architecture of choice in natural language processing, computer vision, speech, and various other domains. Layer normalization, a crucial yet often overlooked component of the transformer architecture, plays an integral role in stabilizing the training process in transformer-based models. Introduced by Ba et al. (2016), layer normalization standardizes the features at each layer, adjusting and scaling the activations to have zero mean and unit variance within a vector. This normalization is performed independently for each hidden vector in contrast to batch normalization (Ioffe and Szegedy, 2015), which relies on statistics (mean and variance) from a batch of data points. Layer normalization is especially effective when working with long sequences of variable lengths. While the benefits of layer normalization, such as improved training speed and better convergence, are well documented (Ba et al., 2016; Xu et al., 2019; Zhang and Sennrich, 2019; Jiang et al., 2024), its specific effects on the hidden vectors within a model and the global properties of the resulting representations remain surprisingly underexplored.
@@
(b) This figure shows the projection of the original vector along the uniform vector, $(\bm{x}.\hat{\bm{1}})\hat{\bm{1}}$, and the remaining component, $\bm{x^{\boldsymbol{1}}_{\perp}}$.
(c) The component of the original vector, after removing the projection along the uniform vector, is kept ( $\bm{x^{\boldsymbol{1}}_{\perp}}$, shown in yellow).
(d) Normalizing perpendicular component to the uniform vector ($\bm{x^{\boldsymbol{1}}_{\perp}}$) to unit norm.
@@
In this paper, we first discuss the global effects of layer normalization on a vector. The conventional explanation of the LayerNorm operation is usually as follows: standardize each vector by subtracting the mean of its elements, divide by the standard deviation. While this is an accurate procedural definition, we ask a more global question: How does the LayerNorm operation transform a vector in representation space?
We present a novel interpretation of LayerNorm and show that it can be understood in three steps: (i) throw away the component of the given vector along $\boldsymbol{1}=[1,1,1,1,\cdots,1]^{T}\in\mathbb{R}^{d}$, (ii) normalize the resultant vector; and (iii) scale the resulting vector by $\sqrt{d}$, where $d$ is the dimensionality of the vector space. This process is illustrated in Figure 2. This operation throws away the information along $\boldsymbol{1}=[1,1,1,1,\cdots 1]^{T}$, which we call the uniform vector, indicating that the information along the uniform vector may not be important.
We then present the property of “irreversibility," where we show that the information lost in the LayerNorm process cannot be recovered. These results naturally lead to discussion about the importance of information along the uniform vector, which is removed by LayerNorm irreversibly, and its comparison with RMSNorm (Zhang and Sennrich, 2019), a variant of LayerNorm that doesn’t remove the component along the uniform vector and is used to train the latest Llama models (Touvron et al., 2023a, b; Dubey et al., 2024). Figure 1 shows the difference between the two normalization methods. To empirically justify removing the components of hidden vectors along the uniform vectors, the hidden representations should have non-trivial components along the uniform vector.
We empirically show that both LayerNorm-based models naturally produce hidden representations orthogonal to the uniform vector during inference time, thus rendering the removal of the uniform vector in LayerNorm inconsequential in practice. This is true even for RMSNorm-based models, where hidden vectors on average operate othogonal to the uniform vector. With these results, we provide both theoretical and empirical justifications for the redundancy of removing the component of hidden representations along the uniform vector, thus supporting the usage of RMSNorm over LayerNorm, which is also more computationally efficient.
@@
    Provide a simple geometric interpretation of the layer normalization process that offers an intuitive explanation of how LayerNorm transforms a vector in representation space.
@@
    Present LayerNorm in action with experiments depicting its importance in stabilizing the residual stream vectors.
@@
    Show that removing the component along the uniform vector is a redundant step in LayerNorm at inference time. These findings advocate for the adoption of RMSNorm over LayerNorm in LLMs.
2 Re-Introducing Layer Normalization
@@
In this paper, we refer to the combination of step-1 and step-2 as the "standardization step", whereas step-3 is referred to as the "scale-and-shift" step. The removal of the mean of components from each component in step-1 is referred to as the "mean subtraction" step in this paper. For an overview of the computations happening within a decoder layer in modern LLMs, we refer the reader to section A.1.
While the original definition of LayerNorm gives us an operational description, it tells us very little about the global properties of the resulting vectors. For example, with the steps described above, we get very little insight into the norm or the orientation of the standardized vector ($\bm{y}$). All we know is that the components of the vector $\bm{y}$ are standardized to have zero mean and unit variance within the vector. But how is this standardized vector oriented compared to the original vector and what are its norms? Such an explanation was also absent in the original formulation of layer normalization (Ba et al., 2016).
Initial explorations into the geometry of layer normalization by Brody et al. (2023) demonstrated that $\bm{y}$ maintains a norm of $\sqrt{d}$ and is oriented orthogonal to the constant vector $\boldsymbol{1}=[1,1,1,1,\cdots,1]^{T}\in\mathbb{R}^{d}$, where $d$ represents the dimensionality of the vector space of hidden vectors. Furthermore, a popular variant of LayerNorm, known as RMSNorm (Zhang and Sennrich, 2019), which omits the “mean subtraction" in the standardization step, has come out as a viable alternative to LayerNorm. In this paper, we expand upon these foundational studies by presenting a novel interpretation of the global effects of LayerNorm on a vector, which we argue provides a more intuitive and informative description of how LayerNorm modifies a hidden vector in representation space. Additionally, we analyze the representations of models using RMSNorm, showing that despite omitting the "mean subtraction step", RMSNorm produces hidden representations with similar orientations.
@@
$\mu=\frac{1}{d}\boldsymbol{1}^{T}\bm{x}$ (1)
where $\boldsymbol{1}=[1,1,1,1,\cdots 1]^{T}$ such that $\boldsymbol{1}\in\mathbb{R}^{d}$, which we refer to as the uniform vector in this paper. Using the scalar mean value calculated above, we define the mean vector as
$\bm{\mu}=\mu\boldsymbol{1}$ (2)
where $\mu$ is a scalar mean as calculated in equation 1. Thus, the mean vector is a vector with each component equal to the mean of the components of vector $\bm{x}$. Using this, we can also rewrite the standard deviation of the components of $\bm{x}$ as:
@@ 2.1 The Uniform Vector and the Mean Vector
The uniform vector or $\boldsymbol{1}=[1,1,1,1,\cdots 1]^{T}$, is called so because all its components are equal or uniform and set to 1. This vector plays a very important role in the formulation and understanding of layer normalization. An important thing to note here is the norm of the uniform vector: $\|\boldsymbol{1}\|_{2}=\sqrt{\boldsymbol{1}^{T}\boldsymbol{1}}=\sqrt{d}$.
The mean vector in the context of LayerNorm is defined in a non-traditional manner. Traditionally, a mean vector is the sum of a few vectors. But in our definition, the mean vector, $\bm{\mu}=\mu\boldsymbol{1}$, is a scaled version of the uniform vector. It is scaled by a value that is the mean or the average of all the components of the hidden vector and is oriented in the direction of the uniform vector. An interesting property of the mean vector is its norm. Let’s define $\theta_{x1}$ as the angle between vectors $\bm{x}$ and the uniform vector. Then taking the L2 norm of the mean vector:
$\begin{split}\|\bm{\mu}\|_{2}&=\|\mu\boldsymbol{1}\|_{2}=\mu\|\boldsymbol{1}\|_{2}=\left(\frac{1}{d}\boldsymbol{1}^{T}\bm{x}\right)\|\boldsymbol{1}\|_{2}\\ &=\frac{1}{d}\|\boldsymbol{1}\|_{2}^{2}\|\bm{x}\|_{2}\cos\theta_{x1}=\|\bm{x}\|_{2}\cos\theta_{x1}\end{split}$ (5)
Here we replace the formula for the mean from equation 1, expand the inner product between the uniform vector and $\bm{x}$ in terms of their norms and the angle between them, and use the fact that $\|\boldsymbol{1}\|_{2}=\sqrt{d}$.
The norm of the mean vector gives us very important insights about what’s going on in the layer normalization process. Equation 5 shows that the norm of the mean vector is nothing but the projection of the underlying vector $\bm{x}$ along the uniform vector. By definition, the mean vector, $\bm{\mu}=\mu\boldsymbol{1}$, is oriented along the direction of the uniform vector. In other words, given that $\theta$ is the angle between vectors $\bm{x}$ and the uniform vector,
$\begin{split}\|\bm{\mu}\|_{2}=\|\bm{x}\|_{2}\cos\theta_{x1}=\frac{\bm{x}\cdot\boldsymbol{1}}{\|\boldsymbol{1}\|_{2}}=\bm{x}\cdot\hat{\boldsymbol{1}}\end{split}$ (6)
@@
Note that $\hat{\bm{1}}$ is the unit vector corresponding to the uniform vector $\bm{1}$. To the best of our knowledge, the geometrical meaning of the mean vector has not been discussed in literature before our work. In the next section, we incorporate this information to explain the LayerNorm process, giving a very intuitive and informative description of the process.
@@ 2.2 Explanation of Layer Normalization
If we go through the steps of layer normalization as discussed in the beginning of section 2, the component-wise subtraction of the mean of all components of a vector can now be written completely in terms of the vector $\bm{x}$. We define $\bm{x^{\boldsymbol{1}}_{\perp}}$ as the component of $\bm{x}$ orthogonal to the uniform vector as follows:
$\bm{x^{\boldsymbol{1}}_{\perp}}=\bm{x}-\bm{\mu}=\bm{x}-(\bm{x}.\hat{\bm{1}})\hat{\bm{1}}$ (8)
@@
$\bm{y}=\sqrt{d}\frac{\bm{x^{\boldsymbol{1}}_{\perp}}}{\|\bm{x^{\boldsymbol{1}}_{\perp}}\|_{2}}$ (9)
where $d$ is the dimensionality of the vector being normalized. This shows that layer normalization can be simply defined as the normalization of the component of a vector orthogonal to the uniform vector, accompanied by a scaling factor. With this, we present the simple and elegant definition of the layer normalization process with a deep geometric meaning. The most intuitive recipe of the layer normalization process is shown in Algorithm 1.
Algorithm 1 : The Standardization Step in Layer Nor
1:Throw away the component of a vector along the uniform vector, $\boldsymbol{1}=[1,1,1,1,\cdots 1]^{T}$
2:Normalize the remaining vector
3:Scale the resulting vector by $\sqrt{d}$, where $d$ is the dimensionality of the vector space
Equation 9 also shows that the norm of the standardized vector is $\sqrt{d}$ and is oriented orthogonal to the direction of the uniform vector, which means it exists in a subspace orthogonal to the uniform vector.
@@ 2.3 The Irreversibility of Layer Normalization
The idea of reversibility is discussed briefly while introducing batch normalization (Ioffe and Szegedy, 2015), which normalizes each feature dimension in the input vector $\bm{x}$ independently by calculating the mean and standard deviation statistics over a large sample of such vectors. Specifically, let $\bm{x}^{1},\bm{x}^{2},\dots\bm{x}^{b}$ be a set of $b$ vectors in the training set, each vector of dimension $d$. Then, both the batch normalization and layer normalization processes for each component can be represented in the following two-step process:
@@
The key difference between batch normalization and layer normalization lies in how the expectation values and variance are calculated for each component. In the case of batch normalization, the mean and variance for each dimension are calculated over the training set. This means that $\mathbb{E}[x_{i}]$ and $Var(x_{i})$ are the same for each vector $\bm{x}^{j}\in\{\bm{x}^{1},\bm{x}^{2},\dots\bm{x}^{b}\}$, but different for each component of the vector. In total, for a $d$-dimensional hidden vector space, there are $2d$ statistics in batch normalization, $d$ for the mean and $d$ for variance. Due to this, the information lost during the standardization step in batch normalization, if important, can be recovered in the scale-and-shift step by simply learning $\alpha_{i}=\sqrt{Var(x_{i})}$ and $\beta_{i}=\mathbb{E}[x_{i}]$, since we have $2d$ learnable parameters in batch normalization. The scale-and-shift step was a very conscious design choice of the inventors of batch normalization to allow for batch normalization to represent an identity transformation if the original hidden representations were optimal for the network.
@@
A subtle but very important difference with layer normalization is how the expectation and variances are calculated for each component. For layer normalization, these statistics are calculated for each vector independently. $\mathbb{E}[x_{i}]$ and $Var(x_{i})$ in layer normalization are the same for each component $x_{i}$ within a vector but are different for different vectors. This means that two variables are created in LayerNorm per hidden vector. As LLMs will normalize way more than $d$ vectors throughout the course of their training, thus creating more than $2d$ statistics, the number of statistic variables used in LayerNorm is very high. Thus, the information lost during layer normalization cannot be recovered by learning just $2d$ components of the $\bm{\alpha}$ and $\bm{\beta}$ learnable parameters. This shows that information once lost during layer normalization cannot be recovered during the scale-and-shift step, making the layer normalization process irreversible.
(a) GPT-J
(b) Pythia 6.9
(c) Llama-3
Model Model Num Num Norm.
Name Dim (d) Params Layers Type